Franklin Templeton has just invested $10,560 for his son (age one). This money will be used for his son’s education 19 years from now. He calculates that he will need $38,123 by the time the boy goes to school.

What rate of return will Mr. Templeton need in order to achieve this goal? Use Appendix B for an approximate answer, but calculate your final answer using the formula and financial calculator methods. (Do not round intermediate calculations. Round your final answer to 2 decimal places.)
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

Hence,

$ 38,123 = $ 10,560* ( 1+r/100 ) ^19

$ 38,123/ $ 10,560 = ( 1+r/100 ) ^19

or (3.610132576) ^ ( 1/19) = 1+r/100

or [{(3.610132576 ) ^ ( 1/19)} -1 ] *100 = r

or 6.99% = r

Hence, the correct answer is 6.99%